Use a Karnaugh map to minimize the POS expression (X+Y)(W +Ž)(X +7 +2)(W+X+Y+Z)

The answer of the question is:

(b) (W + X)(W+Z)(X+Y)(W + X+Y+Z) DYDPlease show me the solution to get this answer. Thanks
Answer #1

Given FlWix, y, z) = (x+T) (W+Ž) In POS X ,7,ZI 1 = (x+y) (wtz) (E+y+z) (W+x+y+2) POS À , ğ, žio is present take i X, Y, Z, wBy using ko map to o minimize minimize the the pos expression Karnaugh xyz et (w+x) (w + Z) slwt +972) (x+7) SO F(WIX,Y,Z) =